Given csc theta = 2 sqrt3/3 and -270 degrees< theta< 180 degrees
a) determine the exact values of theta
b) determine the coordinates of the points on the unit circle where the terminal arm oof theta intersect the unit circle
c)determine the exact values of cos theta and tan theta

Answer by Alan3354(69443)

Given csc theta = 2 sqrt3/3 and -270 degrees< theta< 180 degrees
---
That's Q2
-----
a) determine the exact values of theta
sin = 1/csc = sqrt(3)/2
theta = 120 degs
----------------
b) determine the coordinates of the points on the unit circle where the terminal arm oof theta intersect the unit circle
x = cos(theta)
y = sin(theta)
------
c)determine the exact values of cos theta and tan theta
cos(theta) = -1/2
tan = sin/cos = -sqrt(3)